OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (anonymous):

(5,3)(8,−2) Slope is equal to the change in y over the change in x, or 'rise over run'. m=change in ychange in x The change in x is equal to the difference in x-coordinates (also called run), and the change in y is equal to the difference in y-coordinates (also called rise). m=y2−y1x2−x1 Substitute in the values of x and y into the equation to find the slope. m=−2−(3)8−(5) Simplify. More Detail m=−53

OpenStudy (anonymous):

-5/3